OpenStudy (anonymous):

The time required to finish a test is normally distributed with a mean of 60 minutes and a standard deviation of 10 minutes. What is the z-Score for a student who finishes the test in 45 minutes?

OpenStudy (e.cociuba):

z=45-60/10= -15/10 ANSWER:-1.5

OpenStudy (anonymous):

can you change that to a percentage?